# On average population levels for models with directed diffusion in heterogeneous environments

### v1

@@ Abstract
In 2006 (J. Differential Equ.), Lou proved that, once the intrinsic growth rate $r$ in the logistic model is proportional to the spatially heterogeneous carrying capacity $K$ ($r=K^{1}$), the total population under the regular diffusion exceeds the total of the carrying capacity. He also conjectured that the dependency of the total population on the diffusion coefficient is unimodal, increasing to its maximum and then decreasing to the asymptote which is the total of the carrying capacity. DeAngelis et al (J. Math. Biol. 2016) argued that the prevalence of the population over the carrying capacity is only observed when the growth rate and the carrying capacity are positively correlated, at least for slow dispersal. Guo et al (J. Math. Biol. 2020) justified that, once $r$ is constant ($r=K^{0}$), the total population is less than the cumulative carrying capacity. Our paper fills up the gap for when $r=K^{\lambda}$ for any real $\lambda$, disproving an assumption that there is a critical $\lambda^{\ast}\in(0,1)$ at which the tendency of the prevalence of the carrying capacity over the total population size changes, demonstrating instead that the relationship is more complicated. In addition, we explore the dependency of the total population size on the diffusion coefficient when the third parameter of the dispersal strategy $P$ is involved: the diffusion term is $d\Delta(u/P)$, not just $d\Delta u$, for any $\lambda$. We outline some differences from the random diffusion case, in particular, concerning the profile of the total population as a function of the diffusion coefficient.
@@ 1 Introduction
$\begin{cases}\frac{\partial u}{\partial t}(x,t)=d\Delta u(t,x)+r(x)u(t,x)\left(1-\frac{u(t,x)}{K(x)}\right)=0,&x\in\Omega,~~t>0\\ u(x,0)=u_{0}(x),&x\in\Omega\\ \frac{\partial u}{\partial n}(t,x)=0,&x\in\partial\Omega,~~t>0.\end{cases}$ (1.1)
@@
On the other hand, in [18], the authors justified that when $r(x)$ is constant, and therefore is independent of $K(x)$, the logistic equation with spatially heterogeneous resources always supports a total population strictly smaller than the total carrying capacity at equilibrium, which is just opposite to the case $r=\alpha K$.
@@
Unlike this local approach, in [25] inequality (1.3) was justified globally (and the opposite in [18] for the constant growth rate), for any $d>0$. In addition, it was noticed that the total solution equals the total of the carrying capacity when either $d=0$ or $d\to\infty$ when $r=\alpha K$. Then, a maximum of the average should be attained for some $d\in(0,\infty)$. Moreover, it was conjectured that
@@
is a unimodal function of $d$, increasing from $\int_{\Omega}K~dx$ to its maximum, and then decreasing, approaching the total carrying capacity from above for $d\to\infty$.
@@
  3. 3.
    We demonstrate that, when $r=\alpha\frac{K}{P}$, the limits of $M(d)$ as $d\to 0^{+}$ and $d\to\infty$ are both equal to the integral of the carrying capacity. However, unlike the case of $P\equiv 1$ associated with the random (regular) diffusion, there are examples when $M(d)$ is not unimodal. Moreover, when the limits of $M(d)$ as $d\to 0^{+}$ and $d\to\infty$ are not equal, various forms of $M(d)$ are possible, from monotone to unimodal, or of a more complicated form.
@@ 2 Limit Population Sizes for Very Slow or Very Fast Diffusion
(A)   $r$, $K$, and $P$ are non-constant, positive parameters over $\overline{\Omega}$ such that $r$, $K$, $P\in C^{2}(\overline{\Omega})$.
@@ 3.1 The choice of the dispersal strategy to guarantee abundance
In the case of $P=\alpha\frac{K}{r}$, it follows from Lemma 2.1 and from Corollary 2.3 that $\int_{\Omega}u_{d}\,dx$ converges to $\int_{\Omega}K\,dx$ when both $d\to 0^{+}$ and $d\to+\infty$. And since $\int_{\Omega}u_{d}\,dx>\int_{\Omega}K\,dx$ for every $d>0$, the total population of solutions to (3.1) should be maximized for an intermediate diffusion coefficient $d^{*}\in(0,+\infty)$.
@@
For the case where $P\equiv$ const, Lou [25] conjectures that $M(d)=\int_{\Omega}u_{d}\,dx$ should be a unimodal function. However, our analysis shows that for a general non-constant dispersal strategy $P(x)$, this unimodal behaviour might not hold, as illustrated in Example 4.3.
@@
The proof literally repeats the agrument in [18] for $P\equiv 1$, but it is presented here for completeness.
@@ 3.2 Tendencies for slow dispersal
    If $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx>0$, then for $d>0$ small,
@@
    If $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx<0$ , then for $d>0$ small,
@@
which is negative for small enough $d>0$, since
$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx=-\int_{\Omega}\frac{1}{r^{2}}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx\leq-\frac{1}{\max_{\bar{\Omega}}r^{2}}\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx<0.$
Remark 3.5.
When $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx=0$, nothing can be concluded about the monotonicity of $\int_{\Omega}u_{d}\,dx$ for $d>0$ small, as shown in Example 4.1.
@@
    $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla K\,dx=\int_{\Omega}\left(\frac{|\nabla K|^{2}}{P}-K\frac{\nabla P\cdot\nabla K}{P^{2}}\right)\,dx.$
@@
    $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla K\,dx=\int_{\Omega}\left(P-Kh^{\prime}(K)\right)\frac{|\nabla K|^{2}}{P^{2}}\,dx=\int_{\Omega}\left(h(K)-Kh^{\prime}(K)\right)\frac{|\nabla K|^{2}}{P^{2}}\,dx.$
    If $h^{\prime}(t)>\frac{h(t)}{t}$ for all $t>0$, then $h(t)-th^{\prime}(t)<0$ for all $t>0$, and $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla K\,dx<0$, which implies, by Lemma 3.4, that
@@
    And if $h^{\prime}(t)<\frac{h(t)}{t}$ for all $t>0$, then $h(t)-th^{\prime}(t)>0$, implying $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla K\,dx>0$, and follows from Lemma 3.4 that
@@
$\nabla r=h^{\prime}\left(\frac{K}{P}\right)\nabla\left(\frac{K}{P}\right)$
@@
$\nabla\left(\frac{K}{P}\right)\cdot\nabla r=h^{\prime}\left(\frac{K}{P}\right)\left|\nabla\left(\frac{K}{P}\right)\right|^{2}\geq 0\,\textrm{ everywhere in }\Omega.$
@@
$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx>0.$
@@
When $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx>0$, we can only be sure that $M(d)=\displaystyle\int_{\Omega}u_{d}\,dx$ is maximized at some intermediate dispersion coefficient $d^{*}\in(0,+\infty)$ when $M(0)\geq M(+\infty)$, or equivalently,
@@ 3.3 The case when the correlation is described by a power function
To analyze the total population of $u_{d,\lambda}$ as $d$ increases, let us follow the strategy on Remark 3.9 and analyze the relation between $M_{\lambda}(0)$ and $M_{\lambda}(+\infty)$.
@@ 4 Examples
For small $d>0$, the monotonicity of the total population $M(d)=\int_{\Omega}u_{d}\,dx$ can be determined by Lemma 3.4 by evaluating the sign of $\int_{\Omega}\nabla(K/P)\cdot\nabla r\,dx$. This first example illustrates that, when this integral vanishes, $M(d)$ can be either decreasing or increasing for small values of $d>0$.
Letting $K(x)=(\cos(2\pi x)+2)^{2}$, $P(x)=\sqrt{K(x)}$ and $r(x)=\cos(\pi x)+2$, we obtain
$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx$ $=\int_{0}^{1}(\cos(2\pi x)+2)^{\prime}\cdot(\cos(\pi x)+2)^{\prime}\,dx=\int_{0}^{1}2\pi^{2}\sin(2\pi x)\sin(\pi x)\,dx$
$=\int_{0}^{1}4\pi^{2}\sin^{2}(\pi x)\cos(\pi x)\,dx=\frac{4\pi}{3}\sin^{3}(\pi x)\bigg|_{x=0}^{1}=0.$
And in this case, numerical implementations show that $\int_{\Omega}u_{d}\,dx<\int_{\Omega}K\,dx$ for $d>0$ is small, as seen in Figure 1.
Figure 1: Population decline when $\int_{\Omega}\nabla(K/P)\cdot\nabla r\,dx=0$, for $K(x)=(\cos(2\pi x)+2)^{2}$, $P(x)=\sqrt{K(x)}$, and $r(x)=\cos(\pi x)+2$.
Now, taking $K(x)=(2x^{3}-3x^{2}+3)e^{2x^{3}-3x^{2}}$, $P(x)=e^{2x^{3}-3x^{2}}$ and $r(x)=\cos(2\pi x)+3$, we compute
$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx=\int_{0}^{1}(2x^{3}-3x^{2}+3)^{\prime}\cdot(\cos(2\pi x)+3)^{\prime}\,dx=\int_{0}^{1}-12\pi(x^{2}-x)\sin(2\pi x)\,dx$
$=$ $6(x^{2}-x)\cos(2\pi x)\bigg|_{x=0}^{1}-\int_{0}^{1}6(2x-1)\cos(2\pi x)\,dx$
$=$ $0-\frac{3}{\pi}(2x-1)\sin(2\pi x)\bigg|_{x=0}^{1}+\int_{0}^{1}\frac{3}{\pi}2\sin(2\pi x)\,dx=0-\frac{3}{\pi^{2}}\cos(2\pi x)\bigg|_{x=0}^{1}=0.$
However, in this case, $\int_{\Omega}u_{d}\,dx>\int_{\Omega}K\,dx$ for $d>0$ small, as seen in Figure 2.
Figure 2: Population growth when $\int_{\Omega}\nabla(K/P)\cdot\nabla r\,dx=0$, for $K(x)=(2x^{3}-3x^{2}+3)e^{2x^{3}-3x^{2}}$, $P(x)=e^{2x^{3}-3x^{2}}$ and $r(x)=\cos(2\pi x)+3$.
Therefore, a more refined analysis is needed to determine the monotonicity of $M(d)$ when $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx=0$. For instance, if $r$ is constant on $\overline{\Omega}$, then [18] establishes that $M(d)<\int_{\Omega}K\,dx$ for all $d>0$, which implies that $M(d)$ should be decreasing near $d=0$.
Example 4.2.
This example explores the possible behaviours of $M(d)$ when $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx>0$. We know from Lemma 3.4 that $M(d)$ is increasing near $d=0$, and from Remark 3.9, a global maximum is attained at $d^{*}\in(0,+\infty)$ if $M(0)>M(+\infty)$. This scenario is illustrated in Figure 3a, which was generated using the parameters $K(x)=2+\cos(\pi x)$, $P(x)=1+\frac{1}{5}\cos(\pi x)$ and $r(x)=\frac{5}{4}+\frac{1}{4}\cos(\pi x)$.
@@
Returning to the case where $P$ and $\frac{K}{r}$ are proportional, recall from Remark 3.2 that $M(d)$ exceeds $\int_{\Omega}K\,dx$ for all $d>0$, with $M(d)$ approaching this integral as $d\to 0^{+}$ and $d\to+\infty$. In the case $P\equiv 1$, Lou [25] conjectured that $M(d)$ changes monotonicity only at the point $d^{*}\in(0,+\infty)$ where $M(d)$ attains its maximum.
We now demonstrate that non-constant dispersal strategies $P$ can invalidate this unimodality. In fact, let $K(x)=0.1+\cos(\pi x)+5\cos^{2}(\pi x)-2\cos^{3}(\pi x)$, $P(x)=1.5-3\cos(\pi x)+\cos^{2}(\pi x)+3\cos^{6}(\pi x)$ and $r(x)=\frac{K(x)}{P(x)}$. As shown in Figure 4, the corresponding total population $M(d)$ changes monotonicity three times and exhibits two different local maxima.
@@
Investigations are being conducted to examine the necessary conditions for the parameters $P$ and $K$ for which $M(d)$ changes monotonicity exactly once, which would be a key step toward proving Lou’s conjecture.
@@
For $\lambda=0.5$, the function $M_{\lambda}(d)$ takes a unimodal form, initially increasing due to the positive correlation between $r$ and $\frac{K}{P}$ (Theorem 3.8), reaching a maximum value, and eventually decreasing to a value below $M_{\lambda}(0)$. For $\lambda=1$, following Lou’s conjecture, $M_{\lambda}(d)$ is unimodal with equal limit values as $d\to 0^{+}$ and $d\to+\infty$, thereby validating Theorem 3.1 and Remark 3.2.
@@ 5 Discussion
    In the particular case when the growth rate is a power function of the ratio $r=(K/P)^{\lambda}$, we can characterize the global population size for large disparsal speed. We conclude that, once the species adopt a large enough diffusion rate, the total population exceeds the total of the carrying capacity, once the power function is convex, and is less if the function is concave. This fact has never been established before for the logistic equation with a regular diffusion. The two cases that have been already investigated in the particular case $P\equiv 1$ are $\lambda=1$ [25] and $\lambda=0$ [18]. We illustrate that the values $\lambda=0$ and $\lambda=1$ are in some sense critical.
@@
      Prove or disprove for $P\equiv 1$ the conjecture that for $r$ proportional to $K$, the total population is a unimodal function in $d$.
@@
      If the previous item is justified, what are the conditions on $P,r,K$ (other than constant $P$ and proportional $r,K$) allowing to conclude that the total population is a unimodal function in $d$?
    4. (d)
      What are the relations of $P,r,K$ leading to the monotone dependency of the total population on the diffusion coefficient?
    5. (e)

### v2

@@ Abstract
In 2006 (J. Differential Equ.), Lou proved that, once the intrinsic growth rate $r$ in the logistic model is proportional to the spatially heterogeneous carrying capacity $K$ ($r=K^{1}$), the total population under the regular diffusion exceeds the total of the carrying capacity. DeAngelis et al (J. Math. Biol. 2016) argued that the prevalence of the population over the carrying capacity is only observed when the growth rate and the carrying capacity are positively correlated, at least for slow dispersal. Guo et al (J. Math. Biol. 2020) justified that, once $r$ is constant ($r=K^{0}$), the total population is less than the cumulative carrying capacity. Our paper fills up the gap for when $r=K^{\lambda}$ for any real $\lambda$, disproving an assumption that there is a critical $\lambda^{\ast}\in(0,1)$ at which the tendency of the prevalence of the carrying capacity over the total population size changes, demonstrating instead that the relationship is more complicated. In addition, we explore the dependency of the total population size on the diffusion coefficient when the third parameter of the dispersal strategy $P$ is involved: the diffusion term is $d\Delta(u/P)$, not just $d\Delta u$, for any $\lambda$. We outline some differences from the random diffusion case, in particular, concerning the profile of the total population as a function of the diffusion coefficient.
@@ 1 Introduction
$\begin{cases}\frac{\partial u}{\partial t}(x,t)=d\Delta u(t,x)+r(x)u(t,x)\left(1-\frac{u(t,x)}{K(x)}\right),&x\in\Omega,~~t>0\\ u(x,0)=u_{0}(x),&x\in\Omega\\ \frac{\partial u}{\partial n}(t,x)=0,&x\in\partial\Omega,~~t>0.\end{cases}$ (1.1)
@@
On the other hand, in [18], the authors justified that when $r(x)$ is constant, and therefore is independent of $K(x)$, the logistic equation with spatially heterogeneous resources always supports a total population strictly smaller than the total carrying capacity at equilibrium, which is just opposite to the case $r=\alpha K$. This result can be interpreted as, once the growth rate of species is unchanged over the domain, a uniform distribution of resources maximizes the population of species. To strengthen this result, Mazari-Fouquer [28] quantifies the optimality of a constant carrying capacity on the maximization of average population of species, compared to other resource distribution profiles.
@@
Unlike this local approach, in [26] inequality (1.3) was justified globally (and the opposite in [18] for the constant growth rate), for any $d>0$. In addition, it was noticed that the total solution equals the total of the carrying capacity when either $d=0$ or $d\to\infty$ when $r=\alpha K$. Then, a maximum of the average should be attained for some $d\in(0,\infty)$. Although [26] conjectures that
@@
attains exactly one local maximum in $(0,+\infty)$, it was proven by [25] the existence of a carrying capacity profile that induces multiple local maxima for the average of $u_{d}$.
@@ 2 Limit Population Sizes for Very Slow or Very Fast Diffusion
(A)   $r$, $K$, and $P$ are non-constant, positive functions over $\overline{\Omega}$ such that $r$, $K$, $P\in C^{2}(\overline{\Omega})$.
@@ 3.1 The choice of the dispersal strategy to guarantee abundance
In the case of $P=\alpha\frac{K}{r}$, it follows from Lemma 2.1 and from Corollary 2.3 that $\int_{\Omega}u_{d}\,dx$ converges to $\int_{\Omega}K\,dx$ when both $d\to 0^{+}$ and $d\to+\infty$. And since $\int_{\Omega}u_{d}\,dx>\int_{\Omega}K\,dx$ for every $d>0$, the total population of solutions to (3.1) should be maximized for an intermediate diffusion coefficient $d^{*}\in(0,+\infty)$. However, multiple local maxima for $\int_{\Omega}u_{d}\,dx$ might exist, as illustrated in Example 4.2.
@@
The proof literally repeats the argument in [18] for $P\equiv 1$, but it is presented here for completeness.
@@ 3.2 Tendencies for slow dispersal
    If $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx<0$, then for $d>0$ small,
@@
    If $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx>0$ , then for $d>0$ small,
@@
which is negative for small enough $d>0$, since $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx<0$.
@@
    $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{K}\right)\,dx=\int_{\Omega}\left(-\frac{|\nabla K|^{2}}{PK^{2}}+K\frac{\nabla P\cdot\nabla K}{P^{2}K^{2}}\right)\,dx.$
@@
    $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{K}\right)\,dx=\int_{\Omega}\left(-P+Kh^{\prime}(K)\right)\frac{|\nabla K|^{2}}{P^{2}K^{2}}\,dx=\int_{\Omega}\left(-h(K)+Kh^{\prime}(K)\right)\frac{|\nabla K|^{2}}{P^{2}K^{2}}\,dx.$
    If $h^{\prime}(t)>\frac{h(t)}{t}$ for all $t>0$, then $h(t)+th^{\prime}(t)>0$ for all $t>0$, and $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{K}\right)\,dx>0$, which implies, by Lemma 3.4, that
@@
    And if $h^{\prime}(t)<\frac{h(t)}{t}$ for all $t>0$, then $h(t)+th^{\prime}(t)<0$, implying $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{K}\right)\,dx<0$, and follows from Lemma 3.4 that
@@
$\nabla\left(\frac{1}{r}\right)=-h\left(\frac{K}{P}\right)^{-2}h^{\prime}\left(\frac{K}{P}\right)\nabla\left(\frac{K}{P}\right)$
@@
$\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)=-\frac{h^{\prime}\left(K/P\right)}{h\left(K/P\right)^{2}}\left|\nabla\left(\frac{K}{P}\right)\right|^{2}\leq 0\,\textrm{ everywhere in }\Omega.$
@@
$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx<0.$
@@
When $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx<0$, we can only be sure that $M(d)=\displaystyle\int_{\Omega}u_{d}\,dx$ is maximized at some intermediate dispersion coefficient $d^{*}\in(0,+\infty)$ when $M(0)\geq M(+\infty)$, or equivalently,
@@ 3.3 The case when the correlation is described by a power function
To analyze the total population of $u_{d,\lambda}$ as $d$ increases, let us follow the strategy of Remark 3.8 and analyze the relation between $M_{\lambda}(0)$ and $M_{\lambda}(+\infty)$.
@@ 4 Examples
This example explores the possible behaviours of $M(d)$ when $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx<0$. We know from Lemma 3.4 that $M(d)$ is increasing near $d=0$, and from Remark 3.8, a global maximum is attained at $d^{*}\in(0,+\infty)$ if $M(0)>M(+\infty)$. This scenario is illustrated in Figure 1a, which was generated using the parameters $K(x)=2+\cos(\pi x)$, $P(x)=1+\frac{1}{5}\cos(\pi x)$ and $r(x)=\frac{5}{4}+\frac{1}{4}\cos(\pi x)$.
@@
Returning to the case where $P$ and $\frac{K}{r}$ are proportional, recall from Remark 3.2 that $M(d)$ exceeds $\int_{\Omega}K\,dx$ for all $d>0$, with $M(d)$ approaching this integral as $d\to 0^{+}$ and $d\to+\infty$, implying that $M(d)$ attains a maximum value at $d^{*}\in(0,+\infty)$.
We now illustrate the local maximum value of $M(d)$ my not be unique. Let $K(x)=0.1+\cos(\pi x)+5\cos^{2}(\pi x)-2\cos^{3}(\pi x)$, $P(x)=1.5-3\cos(\pi x)+\cos^{2}(\pi x)+3\cos^{6}(\pi x)$ and $r(x)=\frac{K(x)}{P(x)}$. As shown in Figure 2, the corresponding total population $M(d)$ exhibits two different local maxima.
@@
Investigations are being conducted to examine the necessary conditions for the parameters $P$ and $K$ for which $M(d)$ changes monotonicity exactly once and attains one unique maximum value.
@@
For $\lambda=0.5$, the function $M_{\lambda}(d)$ takes a unimodal form, initially increasing due to the positive correlation between $r$ and $\frac{K}{P}$ (Theorem 3.7), reaching a maximum value, and eventually decreasing to a value below $M_{\lambda}(0)$. For $\lambda=1$, $M_{\lambda}(d)$ is always greater than $\int_{\Omega}K\,dx$ with equal limit values as $d\to 0^{+}$ and $d\to+\infty$, thereby validating Theorem 3.1 and Remark 3.2.
@@ 5 Discussion
    In the particular case when the growth rate is a power function of the ratio $r=(K/P)^{\lambda}$, we can characterize the global population size for large dispersal speed. We conclude that, once the species adopt a large enough diffusion rate, the total population exceeds the total of the carrying capacity, once the power function is convex, and is less if the function is concave. This fact has never been established before for the logistic equation with a regular diffusion. The two cases that have been already investigated in the particular case $P\equiv 1$ are $\lambda=1$ [26] and $\lambda=0$ [18]. We illustrate that the values $\lambda=0$ and $\lambda=1$ are in some sense critical.
@@
      What are the relations of $P,r,K$ leading to the monotone or the unimodal dependency of the total population on the diffusion coefficient?